create 1000 mL of a 5% solution of dextrose using a 50% stock solution and a 1 L bag of 0.9 % saline
Answer #1

C1V1 = C2V2

50%V1 = 5% x 1000

= 50/100 x v1 = 5/100 x 100

V1 = 100ml of 50% Stock

100ml of 50% Stock + 900 ml 0.9 % saline

= 5% solution of dextrose